OpenStudy (anonymous):

which or the following rational number are equal to -17/10? options : -17, -1.7 -34/20 or 0.17 also please answer quickly

OpenStudy (curry):

-.17

OpenStudy (curry):

and -34/20

OpenStudy (curry):

quick enough for u if it then leave medal

OpenStudy (sasogeek):

-1.7

OpenStudy (anonymous):

-.17 is -17/100

OpenStudy (anonymous):

....

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

-17/10 = -1.7 -17/10 = -34/20

OpenStudy (anonymous):

+1 jim thompson

jimthompson5910 (jim_thompson5910):

just combining two posts lol

OpenStudy (anonymous):

those are your answers

OpenStudy (anonymous):

yeah but the first one was wrong

OpenStudy (anonymous):

and the second wan't correct there was more than 1 answer

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

im about 2 give out another question if u get the answer right and quickly i,ll give u a good answer and ull earn a medal ok im about 2 write it